Two blocks are free to slide along the frictionless wooden Lrack shown below. The block of mass m1 4.92 kg is released from uie position shown, at height 5.00 m above the rial part of the Lrack. Protrudiny from its front end is the north pole of a strong magnet, which repels the north Pole of an idenucal maynel ernbedded in the Lack end of the block of mass 9.40 k iniually at rest. Tie uwo Llocks never uouch. Calculate the maximum height to which m1 rises arLer the elastic collision.
Answer #1

speed of m1 when it hits m2

v = sqrt(2gh) = sqrt(2*9.8*5)

= 9.9 m/s

let the speed after collision be v1 and v2

by the conservation of momentum we have

4.92*9.9 + 0 = 4.92*v1 + 9.4v2 -------(1)

velocity of approach = velocity of recess

v2 - v1 = u1 - u2

v2 - v1 = 9.9 ---------(2)

multiply (2) by 4.92 and adding we have

14.32 v2 = 97.416

v2 = 6.8 m/s

v1 = 6.8-9.9 = -3.097 m

so the height reached by m1 after collision = v1^2 /2g

= (3.097^2) / 2*9.8 = 0.49 m